OpenStudy (anonymous):

The memory card on rogers digital camera can hold z pictures (of equal size). after using up 1/4 of the space of the memory card, Roger took 7 more pictures. The card then had 5/7 of its memory remaining. What is the value of Z?

OpenStudy (anonymous):

(1/4)z + 7 = (2/7)z 7z + 196 = 8z 196 = z